OpenStudy (anonymous):

The hypotenuse of a 30°- 60°- 90° right triangle measures 22 cm. What are the lengths of the longer leg and shorter leg?

OpenStudy (cwrw238):

this triangle has sides in ratio 2: 1 : sqrt3 where 2 = hypotenuse

OpenStudy (cwrw238):

so if 2 corresponds to 22 what does 1 and sqrt3 correspond to?

OpenStudy (anonymous):

i have no idea. i missed this lessson. do i use the pythagreon therom?

OpenStudy (cwrw238):

no - just use proportionality hypotenuse/ short side = 2 / 1 = 22/x where x = shortest side giving 2x = 22 so x = 11 11 is length of shortest side in the same for the longest leg: 2 / sqrt 3 = 22 / x solve for x

OpenStudy (anonymous):

thanks

OpenStudy (cwrw238):

yw